# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v2

@@ A.5 Impact of ViT backbone size
Table 13: Linear probing results on the validation set fMoW-RGB, where we vary the size of the ViT encoder ${\mathcal{L}}$ from ViT-B, ViT-L, and ViT-G. “Last 1/Last 4” refers to using the output representation from just the last 1 or the last 4 ViT layers, respectively.
We also test the impact of the ViT backbone for ExPLoRA, varying the architecture for DinoV2 from ViT-B (86M, $\tt{L}=12$ layers, embedding dimension 768), ViT-L (303M parameters, $\tt{L}=24$ layers, embedding dimension 1024), and ViT-G (1100M parameters, $\tt{L}=40$ layers, embedding dimension 1280) for extended pre-training on fMoW-RGB. The ExPLoRA models we compare against are D-[12]-$r64$ for ViT-B, D-[24]-$r64$ for ViT-L, and D-[32]-$r32$ for ViT-G. We unfreeze the 12th, 24th, and 32nd layers for each of ViT-B, ViT-L, and ViT-G, picking these layers by extending the analysis from section 6.3 to ViT-B and ViT-G. We find that the 12th (last layer) for ViT-B and the 32nd (out of 40) layer for ViT-G output representations with low mean eigenvalues compared to other layers, thus presenting good candidates for unfreezing.
@@
Appendix B Training Details
@@ B.1 Pre-Training
B.2 PEFT Fine-Tuning
@@ B.3 Linear Probing
B.4 Multi-Spectral Images
We use the group-channel ViT-L architecture introduced in Cong et al. (2022). We don’t use DinoV2 since there is no such architecture for DinoV2 pre-training. Input images are $13\times 98\times 98$, representing 13 multi-spectral bands. We follow the configuration in Cong et al. (2022) of dropping bands B1, B9, B10, and use the same grouping strategy. When loading MAE weights to the ViT-L encoder, the patch embeddings do not match and so the patch embedding and group channel encodings are trained from scratch. All other configuration details are the same as for M-ExPLoRA in section B.1, except that we use a base learning rate of $4.5\times 10^{-4}$ for pre-training and train for 50 epochs (given the larger dataset size) on 4 NVIDIA RTX A4000 GPUs for 80 hours.
Fine-tuning details are the same as in B.2.
B.5 Downstream datasets
Hyperparameter and training configuration details are the same as in section B.1 if the images are RGB, and the same as in section B.4 if the images have more channels or are temporal.
B.6 Dataset Licenses
@@
Appendix C Environmental Impact
Following Cong et al. (2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table 14.

### v3

@@ B.5 Impact of ViT backbone size
Table 15: Linear probing results on fMoW-RGB (validation), where we vary the size of the ViT encoder ${\mathcal{L}}$ from ViT-B, ViT-L, and ViT-G. “Last 1/Last 4” refers to using the output representation from just the last 1 or the last 4 ViT layers.
We also test the impact of the ViT backbone for ExPLoRA, varying the architecture for DinoV2 from ViT-B (86M, $\tt{L}=12$ layers, embedding dimension 768), ViT-L (303M parameters, $\tt{L}=24$ layers, embedding dimension 1024), and ViT-G (1100M parameters, $\tt{L}=40$ layers, embedding dimension 1280) for extended pre-training on fMoW-RGB. The ExPLoRA models we compare against are D-[12]-$r64$ for ViT-B, D-[24]-$r64$ for ViT-L, and D-[32]-$r32$ for ViT-G. We unfreeze the 12th, 24th, and 32nd layers for each of ViT-B, ViT-L, and ViT-G, picking these layers by extending the analysis from section 7 to ViT-B and ViT-G. We find that the 12th (last layer) for ViT-B and the 32nd (out of 40) layer for ViT-G output representations with low mean eigenvalues compared to other layers, thus presenting good candidates for unfreezing.
@@
B.6 Additional PEFT baselines for MAE
As a continuation of table 1, we include PEFT methods used on MAE weights, which generally underperform compared with DinoV2. For completeness, these results are in table 16.
Model Arch. PEFT
Pre-train
#Params
Fine-tune
#Params
Top 1 Acc.
SatMAE [17] ViT-L LoRA-$r8$ [28] 303.3M 0.8M 76.10
MAE [26] ViT-L LoRA-$r8$ [28] - 0.8M 76.21
MAE [26] ViT-L DVPT-10 [29] - 0.4M 72.35
MAE [26] ViT-L GVPT-100 [70] - 0.4M 70.86
MAE [26] ViT-L SA2VP [49] - 1.1M 73.55
MAE [26] ViT-L AdaLoRA-$r8$ [72] - 1.2M 75.25
MAE [26] ViT-L Adapter+ [59] - 1.4M 74.10
MAE [26] ViT-L Mona [69] - 7.1M 74.76
M-[L]-$r64$ ViT-L LoRA-$r8$ [28] 18.7M 0.8M 76.55
Table 16: MAE+PEFT results on fMoW-RGB validation split (table 1, contd.). “Pre-train #Params" and “Fine-tune #Params" refer to trainable parameters of the ViT encoder required on the new domain (satellite images).
B.7 Results on UDA Benchmarks
As discussed in section A.2, while ExPLoRA is not a traditional unsupervised domain adaptation (UDA) method, it can serve as an initialization for ViT-based UDA approaches. We demonstrate this compatibility below.
Method Arch. Init. plane bcycl bus car horse knife mcycl person plant sktbrd train truck Mean
CDTrans [66] DEiT IN 97.1 90.5 82.4 77.5 96.6 96.1 93.6 88.6 97.9 86.9 90.3 62.8 88.4
PMTrans [75] ViT-B IN-21k 98.9 93.7 84.5 73.3 99.0 98.0 96.2 67.8 94.2 98.4 96.6 49.0 87.5
TVT [67] ViT-B IN-21k 97.1 92.9 85.3 66.4 97.1 97.1 89.3 75.5 95.0 94.7 94.5 55.1 86.7
TVT* [67] ViT-B IN-21k 95.8 85.8 81.9 68.4 95.9 96.2 91.9 70.3 93.8 93.7 92.9 48.5 84.6
TVT [67] ViT-B DinoV2 98.9 88.7 90.3 64.2 99.3 74.5 95.3 66.0 85.3 94.6 97.9 54.6 84.1
TVT [67] ViT-B Ours 97.0 89.9 89.4 73.8 98.0 88.9 94.4 85.9 93.8 94.5 97.7 54.3 88.2
SSRT [60] ViT-B IN-21k 98.9 87.6 89.1 84.8 98.3 98.7 96.3 81.1 94.9 97.9 94.5 43.1 88.8
SSRT [60] ViT-B DinoV2 99.2 88.1 89.9 85.4 98.4 98.9 97.6 84.8 96.2 97.1 95.4 48.3 89.9
SSRT [60] ViT-B Ours 99.4 88.6 91.4 87.9 98.3 99.1 97.1 88.0 95.9 98.1 96.0 51.2 90.9
Table 17: Classification accuracy (%) on VisDA-2017 (validation). Results marked with * use our reproduced results. Using ExPLoRA initialization improves UDA performance compared to standard ImageNet initialization.
VisDA2017
The VisDA2017 dataset (Peng et al., 2017) contains 152,297 training and 55,388 validation images across 12 object classes. The dataset represents a synthetic-to-real domain shift: training images are synthetically rendered 3D models under various lighting conditions, while validation images are sourced from MS-COCO (Lin et al., 2014).
Table 17 shows ExPLoRA’s effectiveness when combined with TVT (Yang et al., 2023) and SSRT (Sun et al., 2022), two state-of-the-art UDA methods. Using ExPLoRA D-[12]-$r64$ (DinoV2-initialized ViT-B with last layer unfrozen and LoRA-r64 elsewhere) pre-trained on both synthetic and real domains, we outperform traditional ImageNet-21k (Deng et al., 2009) initialization by 1-3% while achieving more balanced per-class accuracy. Most notably, when using SSRT, we achieve a new SoTA accuracy of 90.9%, surpassing both ImageNet and DinoV2 initialization by substantial margins ($\uparrow$2.1and $\uparrow$1.0respectively). These results are particularly significant as they show ExPLoRA’s unsupervised initialization can outperform methods that rely on supervised ImageNet-21k pre-training. The benefits of ExPLoRA initialization are also clear with TVT, where performance rises to 88.2% to match recent SoTA methods (Xu et al., 2021; Zhu et al., 2023)– a marked improvement over both its original results ($\uparrow$2.1%) and DinoV2 initialization ($\uparrow$4.1%). These results demonstrate that ExPLoRA serves as a powerful initialization strategy for UDA, effectively bridging domain gaps while remaining computationally efficient.
B.8 Attention Map Visualizations
To aid our analysis in section 7, we visualize attention scores for different ViT blocks across multiple models, including DinoV2, D-[L]-$r64$ (i.e. the last row of table 3), the second and third rows of table 3, MAE, SatMAE, and M-[L]-$r64$. These visualizations are shown in fig. 9 for 3 different images from the validation set of fMoW-RGB. Since our models are trained without registers, we truncate attention scores more than 5 standard deviations away from the mean, thus removing artifact attention scores with unusually high values on background patches (Darcet et al., 2023).
The visualizations in fig. 9 further support the analysis in section 7. For the Dino models, the attention scores of block 9-10 are diffuse and spread around the central object of the image, with quite a few border pixels highlighted. Conversely, the attention scores of the final layers are concentrated more towards the central object. These visualizations further suggest that the middle layers focus on capturing local properties of the images such as texture, while the final layers capture global semantic information such as object-ness. Interestingly, the initial blocks for the Dino models display sparse attention patterns with spikes on seemingly random patches. This might suggest a form of caching to aid the computation of deeper layers that will extract local or global information.
For the MAE models, we see that the original MAE (pre-trained on natural images) seem to highlight more border pixels in the final layers of the ViT. Post extended pre-training with ExPLoRA, the final layers concentrate attention scores on the central object, more closely resembling the patterns of SatMAE (which was fully pre-trained on satellite images). ExPLoRA is thus able to successfully transfer knowledge from its initialized source-domain weights $W_{S}$ to serve as a foundation model $W_{T}^{\ast}$ on the new target domain $T$.
Appendix C Training Details
@@ C.1 Pre-Training
C.2 PEFT Fine-Tuning
@@ C.3 Linear Probing
C.4 Multi-Spectral Images
We use the group-channel ViT-L architecture introduced in (Cong et al., 2022). We don’t use DinoV2 since there is no such architecture for DinoV2 pre-training. Input images are $13\times 98\times 98$, representing 13 multi-spectral bands. We follow the configuration in Cong et al. (2022) of dropping bands B1, B9, B10, and use the same grouping strategy. When loading MAE weights to the ViT-L encoder, the patch embeddings do not match and so the patch embedding and group channel encodings are trained from scratch. All other configuration details are the same as for M-ExPLoRA in section C.1, except that we use a base learning rate of $4.5\times 10^{-4}$ for pre-training and train for 50 epochs (given the larger dataset size) on 4 NVIDIA RTX A4000 GPUs for 80 hours.
Fine-tuning details are the same as in C.2.
C.5 Downstream datasets
Hyperparameter and training configuration details are the same as in section C.1 if the images are RGB, and the same as in section C.4 if the images have more channels or are temporal.
C.6 Dataset Licenses
@@
Appendix D Environmental Impact
Following (Cong et al., 2022), we compare the carbon footprint of pre-training using ExPLoRA with domain-specific solutions such as SatMAE. We use the carbon footprint calculator proposed by Lacoste et al. (2019). Our results are in table 18.